OpenStudy (anonymous):

A water balloon is dropped from a height of 36 ffet. How long does the balloon take to reach the ground? Use the formula H = -16(t)(t) + 36 Can anyone figure out how to do this? Thanks

OpenStudy (anonymous):

What does H represent?

OpenStudy (anonymous):

hight

OpenStudy (anonymous):

If H represents hight, then the balloon reaches the ground when H-=0. That's: \[-16t^2+36=0\] Solve for t, and this is what you'r looking for.

OpenStudy (anonymous):

Can you solve the equation?

OpenStudy (anonymous):

\[-16t^2+36=0 \implies 16t^2=36 \implies t^2=9/4 \implies t=3/2, t=-3/2\] Take only the positive value of t. So, the balloon will take 3/2 (time unite) before reaching the ground.

OpenStudy (anonymous):

thank you so much!

OpenStudy (anonymous):

You're welcome. I hope that made sense to you.

OpenStudy (anonymous):

it does now :)